# Cosmic microwave background constraints on extended dark matter objects

### v1

@@ 1 Introduction
Despite the fact that dark matter (DM) comprises $\sim 85\%$ of the total amount of matter in our Universe, very little is known about its mass: the potential mass range of DM candidates spans over 80 orders of magnitude. At the high end of this spectrum, macroscopic DM candidates, with masses ranging from comparable to large asteroids to large stars, offer an alternative to particle- and wave- like DM. Such celestial objects, formed in the early universe, are so sizeable that they can in principle be detected via their gravitational effects. As such, important constraints have been set using gravitational (micro-)lensing techniques (e.g. Alcock et al. (1998a); Niikura et al. (2019a, b); Green and Kavanagh (2021); Croon et al. (2020a); Bai et al. (2020); Croon et al. (2020b); Crispim Romão and Croon (2024)), and gravitational waves (e.g. Bird et al. (2016); Ali-Haïmoud et al. (2017); Kavanagh et al. (2018); Bertone et al. (2020); Abbott et al. (2019); Chen and Huang (2020); Franciolini (2021); Croon et al. (2023)) on these objects as DM candidates.
@@
In this work, we generalise the accretion formalism further to account for extended dark objects (EDOs) of arbitrary radius and density profiles. In contrast to Bai et al. (2020), which focused on an analytical calculation of the ionisation fraction, our computations are primarily numerical which allows them to be easily adapted to any mass profile. We introduce a scaling technique that significantly speeds up the numerical calculation. The primary purpose of this work is to find the constraints on the fraction of DM that the extended dark objects can comprise. We find stringent constraints for objects over $10^{6}\,\rm R_{\odot}$ or $10^{4}\rm\,AU$ in the mass range above $10^{3}\,\rm M_{\odot}$, complementary to those found in microlensing studies Croon et al. (2020a); Bai et al. (2020); Croon et al. (2020b) and dominating the constraints found by Graham and Ramani (2023) from the dynamical heating of stars in ultrafaint dwarfs in the high mass range.
@@ 2.1.1 Evolution and thermodynamics of baryonic matter
    Taking the hydrostatic approximation, we can take the baryon flow to vanish, such that $v(r)=0$. This is a valid approximation as long as the gravitational interaction is not too intense, placing limits on EDOs with a radius of the same order as their Schwarchild radius.
@@ 3 Ionisation of the cosmic microwave background
$\frac{{\rm d}x_{e}}{{\rm d}z}$ $=C_{r}(z)\frac{\alpha_{\rm B}(T_{\rm M})}{H(z)(1+z)}\left[nx_{e}^{2}+\left(\frac{m_{e}T_{\rm M}}{2\pi}\right)^{3/2}e^{-\frac{E_{\rm I}}{T_{\rm M}}}(1-x_{e})\right].$ (48)
All the necessary definitions are described in what follows. $H(z)$ is the Hubble parameter, where we will take Planck’s value for $H_{0}=73.2$ km/Mpc/s. $\sigma_{\rm T}\simeq 6.65\times 10^{-25}\,{\rm cm}^{2}$ is the Thomson’s constant, $E_{\rm I}=13.2$ eV is the binding energy for the ground state of the hydrogen atom, $T_{\rm cmb}\simeq 2.35\times 10^{-4}\,(1+z)\,{\rm eV}$ is the already mentioned CMB radiation temperature and $n(z)=\rho(z)/m_{\rm p}$ is the baryon density number. Finally, $C_{r}(z)$ is the effective branching ratio (or the so-called Peebles factor), and it describes the probability that excited electrons in a hydrogen atom will reach the ground state. This factor is given by
$C_{r}(z)=\frac{1+K_{H}\Lambda_{H}n(1-x_{e})}{1+K_{H}(\Lambda_{H}+\beta_{H}(T_{\rm M}))n(1-x_{e})}$ (49)
@@
$\beta_{H}(T_{\rm M})$ $=\alpha_{\rm B}(T_{\rm M})\left(\frac{m_{e}T_{\rm M}}{2\pi}\right)^{3/2}e^{3E_{\rm I}/4T_{\rm M}},$ (50)
@@
are characteristic parameters for the ionisation and recombination rates, respectively, with $\alpha=1/137$ being the fine-structure constant.
@@ 3.2 Depositing EDOs in our universe: relative distances and velocities
Taking into account relative velocities is more complicated. As one might expect, an EDO with a high relative speed compared to the speed of sound of baryon fluid, $v_{\rm rel}>c_{s}$, will be more evasive, leading to a smaller total accretion of matter and therefore emitted radiation. To deal with this scenario we will take the approach from Refs.Ali-Haïmoud and Kamionkowski (2017); Bondi (1952); Bondi and Hoyle (1944), where these effects are considered by shifting the speed of sound at infinity by $c_{\infty}\to\sqrt{c_{\infty}^{2}+v_{\rm rel}^{2}}$, decreasing the Bondi radius for the extended dark matter object. We thus expect $v_{\rm rel}$ to be a stochastic variable, and take it to be controlled by a three-dimensional Gaussian linear distribution with a standard deviation $\langle v_{s}^{2}\rangle^{1/2}$. [^3] For large relative speeds, it can be argued that the spherical accretion approximation should break down. Instead, disk-like accretion of baryonic matter should dominate above a certain threshold for $v_{\rm rel}$ Poulin et al. (2017); Serpico et al. (2020). Including this effect would lead to tighter bounds on EDOs. In this work, we remain conservative and focus on this simplified ${}^{,}$ [^4] Additionally, we do not consider the effect of outflows. See Bosch-Ramon and Bellomo (2020) for a discussion about the appropriateness of accretion approximations in light of potential outflows, which may weaken the constraints, particularly for supersonic relative velocities. Following Ref. Ali-Haïmoud and Kamionkowski (2017), the redshift dependence of the standard deviation is given by Dvorkin et al. (2014)
@@ 3.3 Effect on CMB anisotropies
After recombination at redshift $z\sim 1100$, it is assumed that most electrons and protons are combined into hydrogen, leading to an ionisation fraction of $x_{e}\sim 10^{-4}-10^{-3}$. This background neutrality allows the CMB radiation to move freely through space until today, except for the ignition of stars at $z\sim\mathcal{O}(10)$. Therefore, strong constraints can be placed using the state of the universe through the so-called dark ages, as any deviation from neutrality would be a possible source for the absorption of photons, decreasing the visibility of CMB anisotropies. This is where the existence of EDOs comes into play, as their emitted radiation does not just overlay the CMB, but also heats the matter around it, ionizing the neutral hydrogen and locally causing $x_{e}\to 1$. To place constraints, we need to find how this local change in the ionisation fraction affects the global evolution of the universe as studied in the Peebles case B recombination case in Eqs. (47) and (48). As derived in Ref. Ali-Haïmoud and Kamionkowski (2017), the fraction of the power density that is deposited into the background, called the energy deposition rate, $\dot{\rho}_{\rm dep}(t)$, is given by
@@
In Figure 5, we show how the ionisation fraction is modified by the addition of EDOs with different masses, radii, and dark matter fractions. To constrain these objects against their impact on the CMB anisotropies, we would have to implement the modified $x_{e}(z)$ in a Boltzman code and solve for the CMB power spectra, as it was done for PBHs in Refs. Ali-Haïmoud and Kamionkowski (2017); Aloni et al. (2017). However, following Ref. Bai et al. (2020), we note that we may also compare our EDO results to the existing results from PBHs without explicitly repeating this analysis. In this way, we will impose the constrain that $\Delta x_{e}/x_{e}<\mathcal{O}(1)$ at all times, testing particularly for different masses, radii, and dark matter fractions satisfying $\Delta x_{e}(z=50)<10^{-4}$. [^5] It is reasonable to expect that with better observations this threshold will be constrained to smaller values for $\Delta x_{e}(z=50)$. Since the corrections to the recombination equations from Eqs. (64-65) are subdominant, we can generally predict the impact that a change in $\Delta x_{e}(z=50)$ would have on the EDO constrains by just conveniently shifting $f_{\rm DM}$, as they are related by $\delta\Delta x_{e}/\Delta x_{e}=\delta f_{\rm DM}/f_{\rm DM}$.
@@
Figure 6: Constraints on uniform density EDOs for different dark matter fractions as a function of mass vs radius (right) and for different radii as a function of mass vs dark matter fraction (left). The gray regions mark the limitations of our analysis.
Figure 7: Complementarity of constraints on the maximum allowed dark matter fraction as a function of mass, for boson stars and NFW subhaloes of particular sizes. Shown here are constraints from the non-observation of microlensing signals Croon et al. (2020a, b) in the EROS-2 Alcock et al. (1998b), OGLE-IV Niikura et al. (2019b) (note that this does not include the latest results based on a combination of the OGLE-III and OGLE-IV datasets Mroz et al. (2024)), and Subaru-HSC Niikura et al. (2019c) datasets; gravitational wave signals from the inspiral regime of binary mergers assuming binary formation in the early Universe (Croon et al. (2023)), and collisional ionisation of the CMB due to accretion (this work). The gray regions show the limitations of our analysis.
@@ 4 Discussion
The constraints we find are complementary to those found through the non-observation of gravitational microlensing events and binary mergers. In Figure 7 we show these complementary constraints for boson stars and NFW subhaloes together with the CMB constraints derived in this work. As the gravitational wave constraints were derived assuming the inspiral regime, in which the waveform of these object binaries would be known, the objects in the plot have too low of a compactness ($C\lesssim 10^{-3}$) to be observed by the LVK collaboration, as the inspiral regime would have ended at frequencies outside of its observational window. The apparent open window for stellar mass objects with stellar radius can be closed through gravitational heating of dwarf galaxies Graham and Ramani (2023).
@@ Acknowledgements
We thank Joachim Kopp, Vivian Poulin, and Harikrishnan Ramani for useful discussions. DC and SSM are supported by the STFC under Grant No. ST/T001011/1.

### v2

@@ 1 Introduction
Despite the fact that dark matter (DM) comprises $\sim 85\%$ of the total amount of matter in our Universe, very little is known about its mass: the potential mass range of DM candidates spans over 80 orders of magnitude. At the high end of this spectrum, macroscopic DM candidates, with masses ranging from comparable to large asteroids to large stars, offer an alternative to particle- and wave- like DM. Such celestial objects, formed in the early universe, are so sizeable that they can in principle be detected via their gravitational effects. As such, important constraints have been set using gravitational (micro-)lensing techniques (e.g. Alcock et al. (1998a); Niikura et al. (2019a); Zumalacarregui and Seljak (2018); Niikura et al. (2019b); Green and Kavanagh (2021); Fairbairn et al. (2018); Croon et al. (2020a); Bai et al. (2020); Croon et al. (2020b); Crispim Romão and Croon (2024); DeRocco et al. (2023)), and gravitational waves (e.g. Bird et al. (2016); Ali-Haïmoud et al. (2017); Kavanagh et al. (2018); Bertone et al. (2020); Abbott et al. (2019); Chen and Huang (2020); Franciolini (2021); Croon et al. (2023)) on these objects as DM candidates.
@@
In this work, we generalise the accretion formalism further to account for extended dark objects (EDOs) of arbitrary radius and density profiles. In contrast to Bai et al. (2020), which focused on an analytical calculation of the ionisation fraction, our computations are primarily numerical which allows them to be easily adapted to any mass profile. We introduce a scaling technique that significantly speeds up the numerical calculation. The primary purpose of this work is to find the constraints on the fraction of DM that the extended dark objects can comprise. We find stringent constraints for objects over $10^{6}\,\rm R_{\odot}$ or $10^{4}\rm\,AU$ in the mass range above $10^{3}\,\rm M_{\odot}$, complementary to those found in microlensing Fairbairn et al. (2018); Croon et al. (2020a); Bai et al. (2020); Croon et al. (2020b) and projections such as those found in pulsar timing delay studies Dror et al. (2019), and dominating the constraints found by Graham and Ramani (2023, 2024) from the dynamical heating of stars in ultrafaint dwarfs in the high mass range.
@@ 2.1.1 Evolution and thermodynamics of baryonic matter
    Taking the hydrostatic approximation, we can take the baryon flow to vanish, such that $v(r)=0$. This is a valid approximation as long as the gravitational interaction is not too intense, placing limits on EDOs with a radius of the same order as their Schwarzschild radius.
@@ 3 Ionisation of the cosmic microwave background
$\frac{{\rm d}x_{e}}{{\rm d}z}$ $=C_{r}(z)\frac{\alpha_{\rm B}(T_{\rm M})}{H(z)(1+z)}\left[nx_{e}^{2}+\left(\frac{m_{e}T_{\rm cmb}}{2\pi}\right)^{3/2}e^{-\frac{E_{\rm I}}{T_{\rm cmb}}}(1-x_{e})\right].$ (48)
All the necessary definitions are described in what follows. $H(z)$ is the Hubble parameter, where we will take Planck’s value for $H_{0}=67.44$ km/Mpc/s. $\sigma_{\rm T}\simeq 6.65\times 10^{-25}\,{\rm cm}^{2}$ is the Thomson’s constant, $E_{\rm I}=13.2$ eV is the binding energy for the ground state of the hydrogen atom, $T_{\rm cmb}\simeq 2.35\times 10^{-4}\,(1+z)\,{\rm eV}$ is the already mentioned CMB radiation temperature and $n(z)=\rho(z)/m_{\rm p}$ is the baryon density number. Finally, $C_{r}(z)$ is the effective branching ratio (or the so-called Peebles factor), and it describes the probability that excited electrons in a hydrogen atom will reach the ground state. This factor is given by
$C_{r}(z)=\frac{1+K_{H}\Lambda_{H}n(1-x_{e})}{1+K_{H}(\Lambda_{H}+\beta_{\alpha}(T_{\rm M}))n(1-x_{e})}$ (49)
@@
$\beta_{\alpha}(T_{\rm M})$ $=\alpha_{\rm B}(T_{\rm M})\left(\frac{m_{e}T_{\rm cmb}}{2\pi}\right)^{3/2}e^{-E_{\rm I}/4T_{\rm cmb}},$ (50)
@@
are characteristic parameters for the ionisation and recombination rates, respectively, with $\alpha=1/137$ being the fine-structure constant. Considering higher energy levels of hydrogen and other atom species in the universe requires an extension of these equations that will add sub-percentage corrections to these equations Chluba et al. (2007); Ali-Haïmoud and Hirata (2011); Chluba and Thomas (2013, 2011), but these are beyond the error estimates of this work.
@@ 3.2 Depositing EDOs in our universe: relative distances and velocities
Taking into account relative velocities is more complicated. As one might expect, an EDO with a high relative speed compared to the speed of sound of baryon fluid, $v_{\rm rel}>c_{s}$, will be more evasive, leading to a smaller total accretion of matter and therefore emitted radiation. To deal with this scenario we will take the approach from Refs.Ali-Haïmoud and Kamionkowski (2017); Bondi (1952); Bondi and Hoyle (1944), where these effects are considered by shifting the speed of sound at infinity by $c_{\infty}\to\sqrt{c_{\infty}^{2}+v_{\rm rel}^{2}}$, decreasing the Bondi radius for the extended dark matter object. We thus expect $v_{\rm rel}$ to be a stochastic variable, and take it to be controlled by a three-dimensional Gaussian linear distribution with a standard deviation $\langle v_{s}^{2}\rangle^{1/2}$. For large relative speeds, it can be argued that the spherical accretion approximation should break down. Instead, disk-like accretion of baryonic matter should dominate above a certain threshold for $v_{\rm rel}$ Poulin et al. (2017); Serpico et al. (2020). Including this effect would lead to tighter bounds on EDOs. In this work, we remain conservative and focus on this simplified system. [^3] We do not consider the effect of outflows. See Bosch-Ramon and Bellomo (2020) for a discussion about the appropriateness of accretion approximations in light of potential outflows, which may weaken the constraints, particularly for supersonic relative velocities. Following Ref. Ali-Haïmoud and Kamionkowski (2017), the redshift dependence of the standard deviation is given by Dvorkin et al. (2014)
@@ 3.3 Effect on CMB anisotropies
After recombination at redshift $z\sim 1100$, it is assumed that most electrons and protons are combined into hydrogen, leading to an ionisation fraction of $x_{e}\sim 10^{-4}-10^{-3}$. This background neutrality allows the CMB radiation to move freely through space until today, except for the ignition of stars at $z\sim\mathcal{O}(10)$. Therefore, strong constraints can be placed using the state of the universe through the so-called dark ages, as any deviation from neutrality would be a possible source for the absorption of photons, decreasing the visibility of CMB anisotropies. This is where the existence of EDOs comes into play, as their emitted radiation does not just overlay the CMB, but also heats the matter around it, ionizing the neutral hydrogen and locally causing $x_{e}\to 1$. To place constraints, we need to find how this local change in the ionisation fraction affects the global evolution of the universe as studied in the Peebles case B recombination case in Eqs. (47) and (48). As derived in Ref. Ali-Haïmoud and Kamionkowski (2017), the fraction of the power density that is deposited into the background, called the energy deposition rate, $\dot{\rho}_{\rm dep}(t)$, is given by [^4] This analytical approximation uses the fact that, for the considered range of energies, the main process for photon cooling is via inelastic Compton scattering of electrons. Extending this calculation to allow for all interaction channels for photon cooling would add small corrections to our result, increasing the energy deposition rate and, thus, the obtained constraints. We also assume that all these interactions at a certain redshift occur on the spot. Relaxing this approximation is not trivial and requires taking a numerical approach Slatyer et al. (2009); Slatyer (2016).
@@
In Figure 5, we show how the ionisation fraction is modified by the addition of EDOs with different masses, radii, and dark matter fractions. To constrain these objects against their impact on the CMB anisotropies, we would have to implement the modified $x_{e}(z)$ in a Boltzman code and solve for the CMB power spectra, as it was done for PBHs in Refs. Ali-Haïmoud and Kamionkowski (2017); Aloni et al. (2017). However, following Ref. Bai et al. (2020), we note that we may also compare our EDO results to the existing results from PBHs without explicitly repeating this analysis. This is because all accreting massive objects lead to similar imprints on the CMB independently of their shape or nature. Therefore, we can reuse previous constraints if we know how they modify the universe’s ionisation fraction history.
For this, using PBHs constraints from Ref. Ali-Haïmoud et al. (2017), we can compare how different modifications to the ionisation fraction (Figure 12 in Ref. Ali-Haïmoud et al. (2017)) relate to the obtained constraints (Figure 14 in Ref. Ali-Haïmoud et al. (2017)). Moreover, as these modifications are largest for small redshift, as in Figure 5, we focus on the modifications at $z=50$, which gives $\Delta x_{e}(z=50)<10^{-4}$, equivalent to having $\Delta x_{e}/x_{e}<\mathcal{O}(1)$ at all times. [^5] It is reasonable to expect that with better observations this threshold will be constrained to smaller values for $\Delta x_{e}(z=50)$. Since the corrections to the recombination equations from Eqs. (64-65) are subdominant, we can generally predict the impact that a change in $\Delta x_{e}(z=50)$ would have on the EDO constrains by just conveniently shifting $f_{\rm DM}$, as they are related by $\delta\Delta x_{e}/\Delta x_{e}=\delta f_{\rm DM}/f_{\rm DM}$.
@@
Figure 6: Constraints on uniform density EDOs for different dark matter fractions as a function of mass vs radius (right) and for different radii as a function of mass vs dark matter fraction (left). The gray regions mark the limitations of our analysis, where the EDO’s radius is of order of its corresponding Schwarzschild radius.
Figure 7: Complementarity of constraints on the maximum allowed dark matter fraction as a function of mass, for boson stars and NFW subhaloes of particular sizes. Shown here are constraints from the non-observation of microlensing signals Croon et al. (2020a, b) in the EROS-2 Alcock et al. (1998b), OGLE-IV Niikura et al. (2019b) (note that this does not include the latest results based on a combination of the OGLE-III and OGLE-IV datasets Mroz et al. (2024)), and Subaru-HSC Niikura et al. (2019c) datasets; gravitational wave signals from the inspiral regime of binary mergers assuming binary formation in the early Universe (Croon et al. (2023)), and collisional ionisation of the CMB due to accretion (this work). The gray regions show the limitations of our analysis, corresponding to those EDOs whose radius is comparable to their corresponding Schwarzschild radius, or those whose Bondi radius breaks the isolated EDO assumption (Eq. (59)).
@@ 4 Discussion
The constraints we find are complementary to those found through the non-observation of gravitational microlensing events and binary mergers. In Figure 7 we show these complementary constraints for boson stars and NFW subhaloes together with the CMB constraints derived in this work. As the gravitational wave constraints were derived assuming the inspiral regime, in which the waveform of these object binaries would be known, the objects in the plot have too low of a compactness ($C\lesssim 10^{-3}$) to be observed by the LVK collaboration, as the inspiral regime would have ended at frequencies outside of its observational window. The apparent open window for stellar mass objects with stellar radius can be closed through gravitational heating of dwarf galaxies Graham and Ramani (2023, 2024). There is additionally a wide range of already existing PBHs constraints that may similarly apply to EDOs, such as radio and X-ray emission Manshanden et al. (2019).
@@ Acknowledgements
We thank Joachim Kopp, Vivian Poulin, Harikrishnan Ramani, and Jens Chluba for useful discussions. DC and SSM are supported by the STFC under Grant No. ST/T001011/1.